OpenStudy (anonymous):

You are trying to compare the Fahrenheit and Celsius scales and you have two examples: Temperature A is 50 degrees Celsius and 122 degrees Fahrenheit. Temperature B is 100 degrees Celsius and 212 degrees Fahrenheit. What graph models the relationship between the Fahrenheit and Celsius scales? What is an equation of the line in slope-intercept form?

OpenStudy (anonymous):

line between \((50,122)\) and \((100,212)\) slope is \[\frac{212-112}{100-50}=\frac{90}{50}=\frac{9}{5}\]

OpenStudy (anonymous):

equation you get from the point slope formula, use \((y-122)=\frac{9}{5}(x-50)\) and get \[y=\frac{9}{5}x+32\] or if you prefer \[F=\frac{9}{5}C+32\]

OpenStudy (anonymous):

which?

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

kk thanks sooo much

OpenStudy (anonymous):

yw